Q1 a​Saif has to feed Amani. Mom uses Dr Browns feeding system and GoodStart formula. Can he be done in 10 minutes so he can watch his favorite TV   

            program “Suite Life of Zack and Cody” starting in 10 minutes. Draw the network. Show the critical path and Project completion time

Activity

Time

Immediate Predecessor

A. Grab a sterlized bottle

.5 min

---

B. Turn on the perculator and let the water heat

.2 min

A

C. Find formula in refrigerator and pour in bottle

.5 min

A       

D. Pour water, shake and test temperature

1 min

C,B

E. Pickup Amani and get in a comfortable   

     position. Put pillows and toys around.

1 min

D                              

F. Feed Amani

4 min

E

G. Burp Amani

2 min

F

H. Find remote for T. V.

1 min

E

  b.​Given the following network with activities and times estimated in days,
Answer #1

Network diagram is following:

All the paths and their respective duration are listed below:

Path Duration
A-B-D-E-F-G .5+.2+1+1+4+2 = 8.7
A-B-D-E-H .5+.2+1+1+1 = 3.7
A-C-D-E-F-G .5+.5+1+1+4+2 = 9
A-C-D-E-H .5+.5+1+1+1 = 4

Critical path is the path of longest duration

Critical path is : A-C-D-E-F-G

Duration of critical path is 9 minutes.

Completion time is 9 minutes

Therefore, this can be done within 10 minutes